15. Consider a firm that uses only labour and capital. At the present use of labour and capital. the MP of labour is twice the MP of capital, and the price of labour is four times the price of capital. In order to minimize its costs, the firm should a) increase capital and decrease labour.I b) decrease both capital and labour. c) decrease capital and increase labour d) increase both labour and capital. maintain its present factor mix. 2
Answer #1

Equilibrium occurs when MPL/MPK=PL/PK. here

MRTS=(MPL/MPK)=2 and (PL/PK)=4. That is MRTS<PL/PK. By hiring more capital MPk will be decreased because of diminishing return. So that (MPL/MPK) will be increased. Firm should hire more capital until MPL/MPK=PL/PK

Option a is correct.